OpenStudy (anonymous):

lim x-> infinity (3x/1+2x^2)^x

OpenStudy (sirm3d):

the answer is zero

OpenStudy (anonymous):

i dont need answers..

OpenStudy (anonymous):

i need to do the steps.. i have answers too

OpenStudy (sirm3d):

\[y=\lim_{x \rightarrow \infty} \left( \frac{ 3x }{ 1+2x^2 } \right)^x\]

OpenStudy (sirm3d):

apply ln to both sides of the equation, then evaluate the limit at infinity

OpenStudy (anonymous):

that's what i did too

OpenStudy (sirm3d):

\[\ln y=-\infty \]

OpenStudy (anonymous):

why... and since ln(3x / 1+2x^2 ), can't u do ln3x - ln(1+2x^2)

OpenStudy (sirm3d):

\[\lim_{x \rightarrow \infty}\ln \frac{ 3x }{ 1+2x^2 }=-\infty \]

OpenStudy (sirm3d):

splitting the fraction would only cause you trouble

OpenStudy (anonymous):

how do you know that it goes to -ve infinity?

OpenStudy (sirm3d):



OpenStudy (sirm3d):